OpenStudy (anonymous):

what would be the area of a regular polygon with a perimeter of 10 ft and an apothem of 2 ft ?

OpenStudy (ganpat):

Properties of apothems The apothem a can be used to find the area of any regular n-sided polygon of side length s according to the following formula, which also states that the area is equal to the apothem multiplied by half the perimeter since ns = p. A = (pa)/2 This formula can be derived by partitioning the n-sided polygon into n congruent isosceles triangles, and then noting that the apothem is the height of each triangle, and that the area of a triangle equals half the base times the height. An apothem of a regular polygon will always be a radius of the inscribed circle. It is also the minimum distance between any side of the polygon and its center. This property can also be used to easily derive the formula for the area of a circle, because as the number of sides approaches infinity, the regular polygon's area approaches the area of the inscribed circle of radius r = a. A = (pa)/2 = pi * r2.. courtesy Wikipedia... :)